This is considered analytic geometry and is a question I'm not sure where to start with. The teacher did not really cover how to find tangent lines when graphing circles.
Find an equation for the circle with center S (4,-3) that is tangent to the line with equation x - y = 0.
so far, I've been able to determine the x/y intercept at (0,0) based on the equation, and have graphed part of the line. I am not sure how to find the radius or the equation of the circle.

Answer by josgarithmetic(39621)

Center of circle, S (4,-3);
Line tangent to circle is x-y=0.

A radius with one endpoint at S will intersect the line x-y=0 at a right angle, meaning that particular radius is perpendicular to line x-y=0.

Can you use that help to find your way to a solution and answer?